OpenStudy (anonymous):

Determine the solution set of (x + 1)^2 = 25. {-8, 6} {-8, 4} {-6, 4}

OpenStudy (joannablackwelder):

If you square root both sides, you get x+1=+or-5

OpenStudy (joannablackwelder):

Make sense so far?

OpenStudy (anonymous):

ya

OpenStudy (joannablackwelder):

So, now you have 2 equations. x+1=5 and x+1=-5

OpenStudy (joannablackwelder):

Can you solve from there?

OpenStudy (akashdeepdeb):

\[(x+1)^2 = 25\] Take the square root on both sides. \[\sqrt{(x+1)^2} = \pm \sqrt{25} = \pm 5\] Why is it \(\pm\) ? Because both \((+5)^2 ~~and~~ (-5)^2\) give 25. So, \[(x+1) = \pm 5\] \[x = \pm 5 - 1\] Getting this? :)

OpenStudy (anonymous):

4 an -6

OpenStudy (joannablackwelder):

Good job :)

OpenStudy (anonymous):

dude long distance high five

OpenStudy (joannablackwelder):

*high five :)